# Task-Agnostic Machine Learning-Assisted Inference

### v1

@@ 4 ML-assisted False Discovery Rate Control
We introduce PSPS-BH and PSPS-knockoff for ML-assisted FDR control. Briefly, PSPS-BH processes the p-value from ML-assisted linear regression through the BH procedure, while PSPS-knockoff utilizes the ML-assisted debiased Lasso coefficient (Javanmard and Montanari,, 2014; Zhang and Zhang,, 2014; Liu and Rigollet,, 2019), in a ranking algorithm. These methods aim to improve the power of classical methods that employ linear regression or debiased Lasso (Zhang and Zhang,, 2014) using only labeled data. Detailed descriptions of the algorithms are provided below:
Algorithm 2 PSPS-BH for linear regression
Algorithm 3 PSPS-knockoff with debiased Lasso
Here, we employ second-order multivariate Gaussian knockoff variables for knockoff-sample and use the difference between the absolute values of the $k$-th feature and its knockoff coefficient as the knockoff-statistic. Alternative choices for these two steps can also be readily integrated into our algorithm (Ke et al.,, 2024).
We establish the asymptotic FDR control for our proposed method, with the same guarantees as the classical method that only using the labeled data.
Proposition 3.
Assuming positive regression dependence (PRDS) on the p-values (Benjamini and Yekutieli,, 2001), PSPS-BH with ML-assisted linear regression PSPS-LR asymptotically controls the FDR at or below level $qK_{0}/K$, where $K_{0}$ and $K$ is the number of true null and total hypotheses, respectively.
Proposition 4.
PSPS-knockoff with ML-assisted debiased Lasso PSPS-DLasso as the ranking algorithm asymptotically controls the FDR at level $q$.
@@ 5.1 Simulations
We conduct simulations to assess the finite sample performance of our method. Our objectives are to demonstrate that 1) PSPS achieves narrower confidence intervals when applied to statistical tasks already implemented in existing ML-assisted methods; 2) when applied to statistical tasks that have not been implemented for ML-assisted inference, PSPS provides confidence intervals with narrower width but correct coverage (indicating higher statistical power) compared to classical approaches rely solely on labeled data; 3) it provides accurate FDR control and achieves higher power compared to classical methods for FDR control which only use labeled data.
@@
Tasks that have been implemented for ML-assisted inference We compare PSPS with the classical method using only labeled data, the imputation-based method using only unlabeled data, and three ML-assisted inference methods PPI, PPI++, and POP-Inf (Angelopoulos et al., 2023a, ; Angelopoulos et al., 2023b, ; Miao et al., 2023b, ) on mean estimation, linear and logistic. We defer the detailed data generating process to appendix B. In short, we generated outcome $Y_{i}$ from feature $X_{1i}$ and $X_{2i}$, and obtained the pre-trained random forest that predict $Y_{i}$ using $X_{1i}$ and $X_{2i}$. We have 500 labeled samples $(X_{1i},Y_{i},\widehat{f}_{i})$ , and unlabeled samples $(X_{1i},\widehat{f}_{i})$ ranged from 1,000 to 10,000. Our goal is to estimate the mean of $Y_{i}$, the linear and logistic regression coefficient between $Y_{i}$ and $X_{1i}$.
@@
fig. 2a-c shows confidence interval coverage and fig. 2d-f shows confidence interval width. We find that imputation-based method fail to obtain correct coverage, while all other including PSPS have the correct the coverage. PSPS has narrower confidence interval width compared to the classical method and other approaches for ML-assisted inference.
@@
Figure 2: Simulation for implemented tasks including mean estimation, linear and logistic regression from left to right. Panel a-c shows coverage and d-f shows confidence interval width.
Tasks that have not been implemented for ML-assisted inference Next, we consider several commonly used statistical tasks that currently lack implementations for ML-assisted inference including quantile regression (Koenker,, 2005), instrumental variable (IV) regression (Angrist and Imbens,, 1995), negative binomial (NB) regression (Hilbe,, 2011), debiased Lasso (Zhang and Zhang,, 2014), and the Wilcoxon rank-sum test (Mann and Whitney,, 1947). Similar to our previous simulations, we utilize labeled data, unlabeled data, and a pre-trained ML model. Detailed simulation settings are deferred to the appendix B. Our goal is to estimate the regression coefficient between $Y_{i}$ and $X_{1i}$ for quantile (at quantile level 0.75), IV, and NB regression, between $Y_{i}$ and high dimensional features ${\bf X}_{i}\in\mathbb{R}^{150}$ for debiased Lasso, and to perform hypothesis testing on the medians of two independent samples $Y_{i}|X_{1i}=1$ and $Y_{i}|X_{1i}=0$ using the Wilcoxon rank-sum test.
fig. 3a-d shows confidence interval coverage and fig. 3f-i shows confidence interval width for parameter estimation. fig. 3e and fig. 3j shows the type-I error and statistical power for Wilcoxon rank-sum test. We found that imputation-based method fail to obtain correct coverage or have inflated type-I error, while PSPS and classical method have the correct the coverage and type-I error control. PSPS have narrower confidence intervals width in all settings, and higher statistical power for Wilcoxon rank-sum test compared to classical methods. Confidence intervals become narrower as unlabeled sample size increases, indicating higher efficiency gain.
@@
Figure 3: Simulation for not implemented tasks including quantile, instrumental variable (IV), negative binomial (NB) regression, debiased Lasso, and Wilcoxon rank-sum test from left to right. Panel a-e shows coverage (1 - type-error for Wilcoxon rank-sum test) and f-j shows confidence interval width (1 - power for Wilcoxon rank-sum test).
@@
FDR control We evaluate the finite sample performance of PSPS-BH and PSPS-knockoff in controlling the FDR compared with classical and imputation-based methods as baselines. We consider low–dimensional($K<n$) and high-dimensional($K>n$) linear regressions for PSPS-BH and PSPS-knockoff, respectively. We simulate the data such that only a proportion of the features are truly associated with the outcome. The data generating process is deferred to appendix B. Our goal is to select the associated features while maintaining the target FDR level.
@@
fig. 4a-b shows the estimated FDR and fig. 4c-d shows the statistical power for different methods. Imputation-based control method failed to control FDR in either low-dimensional or high-dimensional settings. Classical approach, PSPS-BH, and PSPS-knockoff effectively controlled in both low-dimensional and high-dimensional settings. PSPS-BH, and PSPS-knockoff achieve higher statistical power compared to the classical method.
Figure 4: Simulation for FDR control. Panel a-b shows the estimated FDR level given expected FDR. Panel c-d shows the power.
@@
Together, these simulations demonstrate that PSPS outperforms existing methods and can be easily adopted for a variety of important tasks that have not yet been implemented in existing ML-assisted inference methods.
@@ 5.2 Identify vQTLs for bone mineral density
We employed our method to carry out ML-assisted quantile regression to identify genetic variants associated with the outcome variability (vQTL) of bone mineral density derived from dual-energy X-ray absorptiometry imaging (DXA-BMD) (Miao et al.,, 2022). DXA-BMD serves as the primary diagnostic marker for osteoporosis and fracture risk (Zheng et al.,, 2015; Carey and Delaney,, 2017). Identifying vQTL for DXA-BMD can provide insights into the biological mechanisms underlying outcome plasticity and reveal candidate genetic variants potential gene-gene and gene-environment interactions (Wang et al.,, 2019; Marderstein et al.,, 2021; Westerman et al.,, 2022; Young et al.,, 2018; Miao and Lu,, 2022). We focused on total body DXA-BMD, which integrates measurements from multiple skeletal sites. We used data from the UK Biobank (Bycroft et al.,, 2018), which includes 36,971 labeled and 319,548 unlabeled samples with 9,450,880 genetic variants after quality control. Data processing details have been reported elsewhere (Miao et al.,, 2024). We predicted DXA-BMD values in both labeled and unlabeled samples using SoftImpute (Mazumder et al.,, 2010) with 466 other variables measured in UK Biobank. Prediction in the labeled sample was implemented through cross-validation to avoid overfitting. We used the BH procedure to correct for multiple testing and considered FDR $<0.05$ as the significance threshold.
No genetic variants reached statistical significance under the classical method with only labeled data. PSPS identified 108 significant variants with FDR $<0.05$ spanning 5 independent loci, showcasing the superior statistical power of PSPS (fig. 5 and table 1). Notably, these the significant vQTL cannot be identified by linear regression (Miao et al.,, 2024), indicating the different genetic mechanisms controlling outcome levels and variability for DXA-BMD.
Figure 5: Manhattan plot of vQTLs for bone mineral density. X-axis represents chromosomes (CHR), plotted by base pair positions (BP). Each point on the plot indicates a single nucleotide polymorphism (SNP). Y-axis depicts -log10(p-values).
CHR BP SNP A1 A2 EAF BETA SE p-value FDR
1 205359339 rs12139623 T C 0.105 0.063 0.012 7.2e-08 0.033
4 14359045 rs552582509 A G 0.213 0.046 0.009 4.5e-07 0.045
7 132568586 rs79089873 A G 0.073 -0.068 0.014 5.6e-07 0.049
9 11587905 rs146938822 A G 0.022 -0.134 0.024 4.0e-08 0.033
15 47672201 rs281258 C T 0.393 -0.04 0.007 5.8e-08 0.033
Table 1: Significant vQTLs for bone mineral density. Abbreviations: CHR, Chromosome; BP, Base Pair; SNP, Single Nucleotide Polymorphism; A1, Allele 1 (Effect Allele); A2, Allele 2 (Non-effect Allele); EAF, Effect Allele Frequency; BETA, Effect Size (Beta Coefficient); SE, Standard Error; FDR, False Discovery Rate.
@@ 6 Conclusion
In this work, we introduced a simple, task-agnostic protocol for ML-assisted inference, with applications across a broad range of statistical tasks. We established consistency and asymptotic normality of the proposed estimator. Additionally, we introduce several extensions to further expand the scope of our approach, addressing challenges such as dependence and distributional shifts between labeled and labeled data, and FDR control. Through extensive simulations and applications to real data, we demonstrate the superior performance and broad applicability of our method across diverse tasks. As a future direction, it would be interesting to develop and implement a fast algorithm for resampling in our method.
The codes to implement PSPS, along with the benchmark methods, and to replicate the simulation and real data analysis, are available at https://github.com/qlu-lab/PSPS.
Acknowledgments
The authors gratefully acknowledge research support from National Institutes of Health (NIH) grant U01 HG012039, the University of Wisconsin-Madison Office of the Chancellor, and the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation (WARF).
@@ A.1 Proof the Theorem 1
$n^{1/2}\left(\begin{array}[]{c}\widehat{\bm{\theta}}_{{\cal L}}-\bm{\theta}\\ \widehat{\bm{\eta}}_{{\cal L}}-\bm{\eta}\\ \widehat{\bm{\eta}}_{{\cal U}}-\bm{\eta}\\ \end{array}\right)\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left\{\left(\begin{array}[]{c}{\bf 0}_{K}\\ {\bf 0}_{K}\\ {\bf 0}_{K}\end{array}\right),\left(\begin{array}[]{ccc}{\bf V}(\widehat{\bm{\theta}}_{\cal L})&{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf 0}\\ {\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf V}(\widehat{\bm{\eta}}_{\cal L})&{\bf 0}\\ {\bf 0}&{\bf 0}&\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U})\end{array}\right)\right\},$ (31)
$\widehat{\bm{\theta}}_{{\cal L}}\stackrel{{\scriptstyle P}}{{\rightarrow}}\bm{\theta}$ and $\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bf 0}$. Given fixed weights $\bm{\omega}_{0}=({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$, $\bm{\omega}_{0}^{\rm T}(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bf 0}$.
By Slutsky’s theorem,
$\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}=\widehat{\bm{\theta}}_{{\cal L}}+\bm{\omega}_{0}^{\rm T}(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})\stackrel{{\scriptstyle P}}{{\rightarrow}}\bm{\theta},$ (32)
@@
By multivariate delta methods, denoting $h([{\bf x},\mathbf{y},{\bf z}]^{\rm T})={\bf x}+\bm{\omega}_{0}^{\rm T}({\bf z}-\mathbf{y})$, we have $\nabla h([{\bf x},\mathbf{y},{\bf z}]^{\rm T})=[{\bf 1},-\bm{\omega}_{0},\bm{\omega}_{0}]$, therefore
@@
$\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}(\bm{\theta},{\bf V}(\widehat{\bm{\theta}}_{\cal L})-{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})),$ (37)

### v3

@@ 3.3.3 ML-assisted FDR control
The output from PSPS can be used for ML-assisted FDR control, achieving greater statistical power compared to classical FDR control methods that solely rely on labeled data. We refer to our approach as PSPS-BH and PSPS-knockoff. Briefly, PSPS-BH processes the p-value from ML-assisted linear regression through the Benjamini-Hochberg (BH) procedure [9], while PSPS-knockoff utilizes the ML-assisted debiased Lasso coefficient [23, 51] in the ranking algorithm of knockoff [7]. We present our algorithm in appendix C and evaluate their performance using experiments in section 4.
3.3.4 ML-assisted inference with predicted features
We have discussed ML-assisted inference with outcomes predicted by ML models. Here, we note that PSPS can also be applied when either features alone are predicted or both features and outcomes are predicted. The key idea is that the difference between point estimators obtained from applying $\mathcal{A}$ to predicted features in both labeled and unlabeled datasets is a consistent estimator for zero. This enables zero augmentation for estimators from observed features and outcomes. To implement this, modify step 2 in algorithm 1 to apply $\mathcal{A}$ to predicted features in both labeled and unlabeled data. A similar approach is applicable when both features and outcomes are predicted.
@@ 4.1 Simulations
We conduct simulations to assess the finite sample performance of our method. Our objectives are to demonstrate that 1) PSPS achieves narrower confidence intervals when applied to statistical tasks already implemented in existing ML-assisted methods; 2) when applied to statistical tasks that have not been implemented for ML-assisted inference, PSPS provides confidence intervals with narrower width but correct coverage (indicating higher statistical power) compared to classical approaches rely solely on labeled data; 3) PSPS provides well-calibrated FDR control and achieves higher power compared to classical methods only using labeled data.
@@
Tasks that have been implemented for ML-assisted inference We compare PSPS with the classical method using only labeled data, the imputation-based method using only unlabeled data, and three ML-assisted inference methods PPI, PPI++, and PSPA [3, 4, 35] on mean estimation and linear and logistic regression. We defer the detailed data-generating process to appendix D. In short, we generated outcome $Y_{i}$ from feature $X_{1i}$ and $X_{2i}$, and obtained the pre-trained random forest that predict $Y_{i}$ using $X_{1i}$ and $X_{2i}$. We have 500 labeled samples $(X_{1i},Y_{i},\widehat{f}_{i})$ , and unlabeled samples $(X_{1i},\widehat{f}_{i})$ ranged from 1,000 to 10,000. Our goal is to estimate the mean of $Y_{i}$, as well as the linear and logistic regression coefficient between $Y_{i}$ and $X_{1i}$.
@@
fig. 2a-c show confidence interval coverage and fig. 2d-f show confidence interval width. We find that the imputation-based method fails to obtain correct coverage, while all others including PSPS have the correct coverage. PSPS has narrower confidence intervals compared to the classical method and other approaches for ML-assisted inference.
@@
Figure 2: Simulation for tasks that have been implemented for ML-assisted inference including mean estimation, linear regression, and logistic regression from left to right. Panel a-c present confidence interval coverage and panels d-f present confidence interval width.
@@
Figure 3: Simulation for tasks that have not been implemented for ML-assisted inference including quantile regression, instrumental variable (IV) regression, negative binomial (NB) regression, debiased Lasso, and Wilcoxon rank-sum test from left to right. Panels a-e present confidence interval coverage (1 - type-error for Wilcoxon rank-sum test) and panels f-j present confidence interval width (1 - power for Wilcoxon rank-sum test).
Tasks that have not been implemented for ML-assisted inference Next, we consider several commonly used statistical tasks that currently lack implementations for ML-assisted inference including quantile regression [27], instrumental variable (IV) regression [5], negative binomial (NB) regression [22], debiased Lasso [51], and the Wilcoxon rank-sum test [28]. Similar to our previous simulations, we utilize labeled data, unlabeled data, and a pre-trained ML model. Detailed simulation settings are deferred to the appendix D. Our goal is to estimate the regression coefficient between $Y_{i}$ and $X_{1i}$ for quantile (at quantile level 0.75), IV, and NB regression, between $Y_{i}$ and high dimensional features ${\bf X}_{i}\in\mathbb{R}^{150}$ for debiased Lasso, and to perform hypothesis testing on the medians of two independent samples $Y_{i}|X_{1i}=1$ and $Y_{i}|X_{1i}=0$ using the Wilcoxon rank-sum test.
fig. 3a-d show confidence interval coverage and fig. 3f-i show confidence interval width for parameter estimation. fig. 3e and fig. 3j show the type-I error and statistical power for the Wilcoxon rank-sum test. We found that the imputation-based method fails to obtain correct confidence interval coverage and shows inflated type-I error, while PSPS and classical method have the correct coverage and well-calibrated type-I error control. PSPS has narrower confidence intervals width in all settings, and higher statistical power for the Wilcoxon rank-sum test compared to classical methods. Confidence intervals become narrower as unlabeled sample size increases, indicating higher efficiency gain.
@@
FDR control We evaluate the finite sample performance of PSPS-BH and PSPS-knockoff in controlling the FDR compared with classical and imputation-based methods as baselines. We consider low–dimensional($K<n$) and high-dimensional($K>n$) linear regressions for PSPS-BH and PSPS-knockoff, respectively. We simulate the data such that only a proportion of the features are truly associated with the outcome. The data generating process is deferred to appendix D. Our goal is to select the associated features while maintaining the target FDR level.
@@
fig. E.1a-b shows the estimated FDR and fig. E.1c-d shows the statistical power for different methods. Imputation-based method failed to control FDR in either low-dimensional or high-dimensional settings. Classical approach, PSPS-BH, and PSPS-knockoff effectively controlled in both low-dimensional and high-dimensional settings. PSPS-BH, and PSPS-knockoff achieve higher statistical power compared to the classical method.
@@
These simulations demonstrate that PSPS outperforms existing methods and can be easily adapted for various statistical tasks not yet implemented in current ML-assisted inference methods.
@@ 4.2 Identify vQTLs for bone mineral density
We employed our method to carry out ML-assisted quantile regression to identify genetic variants associated with the outcome variability (vQTL) of bone mineral density derived from dual-energy X-ray absorptiometry imaging (DXA-BMD) [33]. DXA-BMD is the primary diagnostic marker for osteoporosis and fracture risk [54, 15]. Identifying vQTL for DXA-BMD can provide insights into the biological mechanisms underlying outcome plasticity and reveal candidate genetic variants involved in potential gene-gene and gene-environment interactions [45, 29, 47, 49, 32]. We focused on total body DXA-BMD, which integrates measurements from multiple skeletal sites. We used data from the UK Biobank [13], which includes 36,971 labeled and 319,548 unlabeled samples with 9,450,880 genetic variants after quality control. We predicted DXA-BMD values in both labeled and unlabeled samples using SoftImpute [30] with 466 other variables measured in the UK Biobank. Prediction in the labeled sample was implemented through cross-validation to avoid overfitting. The implementation detail is deferred to appendix D. We used the BH procedure to correct for multiple testing and considered FDR $<0.05$ as the significance threshold.
No genetic variants reached statistical significance under the classical method with only labeled data. PSPS identified 108 significant variants with FDR $<0.05$ spanning 5 independent loci, showcasing the superior statistical power of PSPS (fig. E.2 and table E.1). Notably, these significant vQTL cannot be identified by linear regression [36], indicating the different genetic mechanisms controlling outcome levels and variability for DXA-BMD.
4.3 Computational efficiency
We compared the computational efficiency of PSPS with existing methods using a dataset of 500 labeled and 10,000 unlabeled data points. Results are shown in table E.2. While PSPS is slower due to resampling, its overall runtime is still relatively short.
@@ 5 Conclusion
We introduced a simple, task-agnostic protocol for ML-assisted inference, with applications across a broad range of statistical tasks. We established the consistency and asymptotic normality of the proposed estimator. We further introduced several extensions to expand the scope of our approach. Through extensive experiments, we demonstrated the superior performance and broad applicability of our method across diverse tasks. Our protocol involves initially generating summary statistics using computationally efficient software tools in scientific data analysis, followed by integration of summary statistics to produce ML-assisted inference results, which achieves high computational efficiency while maintaining statistical validity. Future work could focus on developing a fast resampling algorithm to further improve computational efficiency.
Acknowledgements: We gratefully acknowledge research support from the National Institutes of Health (NIH; grant U01 HG012039) and support from the University of Wisconsin–Madison Office of the Chancellor and the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation.
@@ A.1 Proof the theorem 1
$\widehat{{\bm{\theta}}}_{{\cal L}}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\theta}}^{*}$ and $\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bf 0}$. Given weights $\widehat{\bm{\omega}}_{0}=(\widehat{\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho\widehat{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}\widehat{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$, where the variances are consistently estimated, Slutsky’s theorem implies that, $\widehat{\bm{\omega}}_{0}^{\rm T}(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bf 0}$.
Also by Slutsky’s theorem,
$\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}=\widehat{{\bm{\theta}}}_{{\cal L}}+\widehat{\bm{\omega}}_{0}^{\rm T}(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\theta}}^{*},$ (21)
@@
By multivariate delta methods, denoting $h([{\bf x},\mathbf{y},{\bf z}]^{\rm T})={\bf x}+\bm{\omega}_{0}^{\rm T}({\bf z}-\mathbf{y})$, we have $\nabla h([{\bf x},\mathbf{y},{\bf z}]^{\rm T})=[{\bf 1},-\bm{\omega}_{0},\bm{\omega}_{0}]$, therefore by the consistency of $\widehat{\bm{\omega}}_{0}$,
@@
$\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bm{\theta}}^{*},{\bf V}(\widehat{\bm{\theta}}_{\cal L})-{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})),$ (26)